Calculate the probability for an electron to be found at an energy of (EF + 2kBT)
in a metal
1
Expert's answer
2020-02-10T09:33:29-0500

Electrons are particles with half-integer spin. And they obey Fermi-Dirac statistics. The function of Fermi-Dirac distribution is written as follows: 

"P(E,T)=\dfrac{1}{exp( \dfrac{E-E_F}{kT})}"

where "P(E,T)" - the probability that the electron occupies an energy level "E" , above or below the Fermi level "E_F"   

Then 

"P(E_F+2k_BT,T)= \dfrac{1}{1+exp(\dfrac{E_F+2k_BT-E_F}{k_BT})}=\dfrac{1}{e^2}=0.119 \newline
in \ percentage \ P(E_F+2k_BT,T) =11.9\%"